# Through the Judge’s Eyes: Inferred Thinking Traces Improve Reliability of LLM Raters

### v1

@@ Abstract.
Large language models (LLMs) are increasingly used as raters for evaluation tasks. However, their reliability is often limited for subjective tasks, when human judgments involve subtle reasoning beyond annotation labels. Thinking traces, the reasoning behind a judgment, are highly informative but challenging to collect and curate. We present a human-LLM collaborative framework to infer thinking traces from label-only annotations. The proposed framework uses a simple and effective rejection sampling method to reconstruct these traces at scale. These inferred thinking traces are applied to two complementary tasks: (1) fine-tuning open LLM raters; and (2) synthesizing clearer annotation guidelines for proprietary LLM raters. Across multiple datasets, our methods lead to significantly improved LLM-human agreement. Additionally, the refined annotation guidelines increase agreement among different LLM models. These results suggest that LLMs can serve as practical proxies for otherwise unrevealed human thinking traces, enabling label-only corpora to be extended into thinking–trace–augmented resources that enhance the reliability of LLM raters. [^1] For reproducing results, our codebase can be accessed at: https://github.com/xingjian-zhang/thru_judge_eye.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
@@
Large language models (LLMs) are increasingly used as automated evaluators for open-ended text generation tasks, primarily due to their remarkable efficiency and scalability (Gu et al., 2024a; Zheng et al., 2023). This approach has been applied to evaluate diverse applications such as code generation (Zhou et al., 2025), text summarization (Bedemariam et al., 2025), and dialogue (Gu et al., 2024a). However, the reliability of LLM evaluators often decreases in subjective tasks that require nuanced human judgment (Ismayilzada et al., 2024; Gómez-Rodríguez and Williams, 2023), especially when they did not receive specialized training (Krumdick et al., 2025).
@@
To understand what is missing in calibrating these models, we can draw an analogy from the process of training human annotators for qualitative coding tasks. To ensure alignment on a subjective task, a novice annotator typically benefits from three key components that an expert can provide: (1) an annotation codebook with detailed instructions and rubrics; (2) a set of examples illustrating representative inputs and their corresponding labels; and critically, (3) explanations detailing the reasoning for assigning a specific label to each example. Among the three components, explanations are particularly important in creating a shared understanding of the task among annotators (Artstein and Poesio, 2008; O’Connor and Joffe, 2020). We hypothesize that these explanations of the expert’s reasoning, which we term the thinking traces, are equally necessary to calibrate LLM evaluators and improve their alignment with human judgment.
@@
Figure 1. Original Annotation Codebook (Evaluating Complexity of Short Stories) (Chhun et al., 2022). Only very basic instructions and scoring rubrics are provided.
To tackle this question, we present a human-LLM collaborative framework to infer thinking traces from label-only human annotations. We employ a simple, yet effective, rejection sampling method to reconstruct these traces at scale. We demonstrate the utility of inferred thinking traces through two complementary applications. First, we use them as training data to fine-tune open-weight LLM raters so that they are specialized for the evaluation task. Second, for models where fine-tuning is not an option, we introduce a novel method for codebook refinement. Inspired by iterative refinement practices in qualitative coding (O’Connor and Joffe, 2020), this approach addresses the common problem of ambiguous instructions in many annotation codebooks (Klie et al., 2023) (Figure 1). Our automatic pipeline leverages the inferred traces to synthesize a new, more explicit codebook that better steers the model’s judgments. Experiments show that both methods significantly improve the alignment of LLM raters with human judgments. Furthermore, the synthesized codebook also improves the agreement between different LLM providers, demonstrating its utility in promoting cross-model consistency. These findings demonstrate that model-inferred thinking traces can effectively serve as a practical proxy for human reasoning, thus improving the reliability and consistency of automated LLM raters on subjective evaluation tasks.
Figure 2. An example of the inferred thinking trace in evaluating the engagement level of a short story.
2. Related Work
@@
Reasoning Language Models (RLMs) represent a significant evolution from standard LLMs, as they are specifically trained to solve tasks that require multiple steps of deliberation. This class of models has become the de facto standard for high performance (Comanici et al., 2025; Guo et al., 2025; Jaech et al., 2024), demonstrating superior capabilities in tasks such as math and coding. Models such as DeepSeek R1 demonstrate the effectiveness of large-scale reinforcement learning in cultivating complex reasoning behaviors in these models (Guo et al., 2025), and provide public access to these reasoning tokens. Although these models are intended to align with “human priors” (Guo et al., 2025), the validity of these reasoning tokens remains a debatable question (Amirizaniani et al., 2024; Dasgupta et al., 2024; Mondorf and Plank, 2024). Testing the validity of thinking traces is beyond the scope of this paper. However, we are interested in a practical objective: leveraging RLMs as a tool to infer latent thinking traces from human experts, thereby improving the reliability of downstream LLM raters.
@@
A powerful paradigm for generating synthetic data is rejection sampling fine-tuning, where a model generates numerous candidate reasoning traces, and only those passing a filtering criterion are used for subsequent training. Existing work has pioneered this approach for tasks with objectively correct answers, such as mathematical reasoning, using an automatic verifier to accept only correct outputs (Guo et al., 2025; Zelikman et al., 2022; Wadhwa et al., 2024; Tong et al., 2024; Singh et al., 2023). To handle more nuanced tasks, methods such as RAFT (Dong et al., 2023) replaced the binary verifier with a learned reward model, filtering for outputs that score highly on desired attributes. This trend of applying reasoning to more subjective domains is also evident in recent work on recommender systems (Tsai et al., 2024). Crucially, while previous work exclusively focuses on using these curated traces for fine-tuning, we introduce a novel second application: distilling insights from the thinking traces to automatically refine annotation codebooks.
@@
The refinement of codebooks is a long-standing challenge in both qualitative research and large-scale data annotation. Research in HCI and crowd-sourcing communities developed human-centric workflows to improve task instructions, often relying on crowd workers to identify ambiguities that were then resolved by an expert or through discussion (Chang et al., 2017; Manam and Quinn, 2018; K. Chaithanya Manam et al., 2019; Pradhan et al., 2022). More recently, researchers have begun to leverage LLMs to automate this process. This includes work on human-LLM collaboration for qualitative coding (Xiao et al., 2023; Halterman and Keith, 2025; Wiebe et al., 2025; Torii et al., 2024; Meng et al., 2024). Other work has demonstrated that LLMs can automatically analyze data or synthesize new guidelines to improve downstream performance (Bibal et al., 2025; Hsu et al., 2025; Srivastava et al., 2025). A key distinction of these automated methods is that they rely only on an LLM’s intrinsic capabilities to analyze content or generate text, rather than being grounded in direct human annotation feedback. In contrast, our approach creates a collaborative pipeline with inputs from both humans and LLMs.
3. Method
3.1. Inferring Latent Thinking Traces
Figure 3. Illustration of Inferring Thinking Traces through an RLM. Details are provided in Section 3.1.
@@ 3.1. Inferring Latent Thinking Traces
Our goal is to generate a high-quality proxy for this latent trace, which we term the reconstructed thinking trace, $t^{\prime}_{i}$. To achieve this, we use an RLM that can generate thinking tokens, hereafter referred to as the generator model, to perform rejection sampling. For each input item $x_{i}$, we provide the generator model with the codebook $\mathcal{C}$} and the annotation target and prompt it to generate a step-by-step thinking trace that concludes with a final rating. This generation is performed $k$ times independently, yielding a set of $k$ candidate trace-label pairs, $\{(t^{\prime}_{i,j},y^{\prime}_{i,j})\}_{j=1}^{k}$ where $t^{\prime}_{i,j}$ denotes the $j$-th candidate trace and $y^{\prime}_{i,j}$ denotes the $j$-th predicted label for $x_{i}$.
For a model to reach the correct human label $y_{i}$, its reasoning is more likely (but not guaranteed) to be a plausible approximation of the latent human thought process. Therefore, we filter the candidate set by retaining only those traces where the model’s predicted label $y^{\prime}_{i,j}$ matches the ground-truth human label $y_{i}$. This creates a set of traces for each sample $x_{i}$:
@@
3.2. Improving LLM Raters
@@
(1) $\mathcal{L}_{\text{SFT}}=\mathbb{E}_{(x,t^{\prime},y)\sim\mathcal{D}_{\text{reason}}}\left[-\log\pi_{\theta}(t^{\prime},y|x)\right]$
@@
We provide the complete prompts for this process in Appendix C.
@@
Step-by-Step Rating:
@@
(1) The story is very simple, lacks depth, and introduces no intricate or developed elements. (2) The story is somewhat simple, with few elements that are underdeveloped, fragmented, or disconnected. (3) The story demonstrates some complexity, introducing multiple elements, but lacks depth, development, or coherence in integrating these elements. (4) The story is mostly complex, weaving together several developed elements with minor areas that could be expanded or refined. (5) The story is highly elaborate and complex, integrating multiple layers (e.g., detailed world-building, character depth, advanced structure, thematic sophistication) in a cohesive and interconnected way.
@@
Figure 4. Refined Annotation Codebook. Part of the content is omitted due to space limitations. See Appendix D for a complete example.
3.3. Evaluation
@@
This criterion measures the degree to which an LLM rater’s outputs agree with human judgments. We apply this evaluation to the outputs of both our applications. For this analysis, we treat a held-out set of human ratings as the gold standard and compare the LLM-generated labels against them. To quantify this agreement, we use several standard metrics for labels in the Likert scale. Our primary focus is on Kendall’s Tau ($\tau$), as ranking correlation [^3] We observe a high correlation between Spearman’s $\rho$ and Kendall’s $\tau$ in the results. We only report $\tau$ here for simplicity. is the most used measure in the literature (Liu et al., 2023, 2024; Thakur et al., 2025; Gu et al., 2024b). For a more complete assessment, we also report the Mean Squared Error (MSE) for average error magnitude and the Intraclass Correlation Coefficient (ICC3) for consistency, based on a two-way mixed-effects model (Koo and Li, 2016).
@@
For codebook refinement, another robust indicator of success is whether $\mathcal{C}$$^{\prime}$ enables independent raters to arrive at the same conclusions consistently. High inter-rater reliability is a sign of a clear and well-defined annotation process. To measure this, we use the improved codebook $\mathcal{C}$$^{\prime}$ to prompt $M$ different proprietary LLMs to rate the same set of items. We then calculate the reliability across these $M$ independent LLM raters using ICC3 (Koo and Li, 2016).
@@
4. Experiment
Task Original Model SFT (DeepSeek Trace) SFT (gpt-oss Trace)
Metric ICC3 $\tau$ MSE ICC3 $\tau$ MSE ICC3 $\tau$ MSE
Short Story (Complexity) 0.231 0.177 2.016 0.604* 0.463* 0.838* 0.581* 0.460* 0.944*
Short Story (Engagement) 0.289 0.275 1.939 0.364 0.048 1.364 0.371 0.179 1.831
Student-Written Essay 0.233 0.202 1.947 0.267 0.322* 1.149* 0.244 0.240 1.387*
News Summary (Consistency) 0.186 0.141 2.904 0.319 0.276 3.476 0.331 0.243 2.970
News Summary (Fluency) 0.378 0.275 1.938 0.402 0.319 1.160* 0.498 0.306 1.200*
Translation 0.240 0.171 2.778 0.348* 0.260* 2.492 0.314* 0.256* 2.439*
Chatbot Response (Correctness) 0.288 0.162 2.049 0.403* 0.297* 1.815* 0.433* 0.263* 1.897
Chatbot Response (Helpfulness) 0.274 0.176 1.969 0.422* 0.264* 1.723* 0.466* 0.295* 1.575*
Average 0.265 0.197 2.193 0.391 0.281 1.752 0.405 0.280 1.780
Table 1. Effect of Reasoning-Enhanced SFT on Human-LLM Rater Agreement. For the metrics, higher is better for ICC3 and Kendall’s $\tau$, while lower is better for MSE. An asterisk (*) denotes a statistically significant improvement over the baseline ($p<0.05$).
4.1. Datasets
@@
  3. (3)
    Evaluating News Summaries. We use the SummEval benchmark (Fabbri et al., 2020), which provides human evaluations of machine-generated summaries of news articles. Each summary is annotated by three expert annotators with prior experience in writing research paper summaries for academic conferences. We focus on the consistency and fluency, which respectively evaluate factual correctness and linguistic quality, two key aspects of summary reliability and readability.
  4. (4)
@@
More statistics and pre-processing details for each dataset are provided in Appendix A.
4.2. Data Processing
@@
As outlined in Section 3, we use rejection sampling to infer thinking traces from the annotation datasets. The original annotation codebook from each source serves directly as the prompt. For our generator models, we selected DeepSeek-R1 and gpt-oss-120b as they output complete thinking traces. We prompt each model to generate sixteen ($k=16$) thinking traces for every sample and keep the first trace that aligns with the human annotator’s final rating. A sampling temperature of 1.0 is used during inference to encourage generative diversity.
4.3. Fine-Tuning Specialized LLM Rater

### v2

@@ Abstract
Large language models (LLMs) are increasingly used as raters for evaluation tasks. However, their reliability is often limited for subjective tasks, when human judgments involve subtle reasoning beyond annotation labels. Thinking traces, the reasoning behind a judgment, are highly informative but challenging to collect and curate. We present a human-LLM collaborative framework to infer thinking traces from label-only annotations. The proposed framework uses a simple and effective rejection sampling method to reconstruct these traces at scale. These inferred thinking traces are applied to two complementary tasks: (1) fine-tuning open LLM raters; and (2) synthesizing clearer annotation guidelines for proprietary LLM raters. Across multiple datasets, our methods lead to significantly improved LLM-human agreement. Additionally, the refined annotation guidelines increase agreement among different LLM models. These results suggest that LLMs can serve as practical proxies for otherwise unrevealed human thinking traces, enabling label-only corpora to be extended into thinking–trace–augmented resources that enhance the reliability of LLM raters. [^1] For reproducing results, our codebase can be accessed at: https://github.com/xingjian-zhang/thru_judge_eye/tree/main
1 Introduction
@@
Large language models (LLMs) are increasingly used as automated evaluators for open-ended text generation tasks, primarily due to their remarkable efficiency and scalability (Gu et al., 2024; Zheng et al., 2023). This approach has been applied to evaluate diverse applications such as code generation (Zhou et al., 2025), text summarization (Bedemariam et al., 2025), and dialogue (Gu et al., 2024). However, the reliability of LLM evaluators often decreases in subjective tasks that require nuanced human judgment (Ismayilzada et al., 2024; Gómez-Rodríguez & Williams, 2023), especially when they did not receive specialized training (Krumdick et al., 2025).
@@
To understand what is missing in calibrating these models, we can draw an analogy from the process of training human annotators for qualitative coding tasks. To ensure alignment on a subjective task, a novice annotator typically benefits from three key components that an expert can provide: (1) an annotation codebook with detailed instructions and rubrics; (2) a set of examples illustrating representative inputs and their corresponding labels; and critically, (3) explanations detailing the reasoning for assigning a specific label to each example. Among the three components, explanations are particularly important in creating a shared understanding of the task among annotators (Artstein & Poesio, 2008; O’Connor & Joffe, 2020). We hypothesize that these explanations of the expert’s reasoning, which we term the thinking traces, are equally necessary to calibrate LLM evaluators and improve their alignment with human judgment.
@@
Figure 1: An example of the annotation codebook for evaluating complexity of short stories (Chhun et al., 2022). Only basic instructions and vague scoring rubrics are provided.
To tackle this question, we present a human-LLM collaborative framework to infer thinking traces from label-only human annotations. We employ a simple, yet effective, rejection sampling method to reconstruct these traces at scale. We demonstrate the utility of inferred thinking traces through two complementary applications. First, we use them as training data to fine-tune open-weight LLM raters so that they are specialized for the evaluation task. Second, for models where fine-tuning is not an option, we introduce a novel method for codebook refinement. Inspired by iterative refinement practices in qualitative coding (O’Connor & Joffe, 2020), this approach addresses the common problem of ambiguous instructions in many annotation codebooks (Klie et al., 2024) (Figure 1). Our automatic pipeline leverages the inferred traces to synthesize a new, more explicit codebook that better steers the model’s judgments. Experiments show that both methods significantly improve the alignment of LLM raters with human judgments. Furthermore, the synthesized codebook also improves the agreement between different LLM providers, demonstrating its utility in promoting cross-model consistency. These findings demonstrate that model-inferred thinking traces can effectively serve as a practical proxy for human reasoning, thus improving the reliability and consistency of automated LLM raters on subjective evaluation tasks.
2 Related Work
@@
Reasoning Language Models (RLMs) represent a significant evolution from standard LLMs, as they are specifically trained to solve tasks that require multiple steps of deliberation. This class of models has become the de facto standard for high performance (Comanici et al., 2025; Guo et al., 2025; Jaech et al., 2024), demonstrating superior capabilities in tasks such as math and coding. Models such as DeepSeek R1 demonstrate the effectiveness of large-scale reinforcement learning in cultivating complex reasoning behaviors in these models (Guo et al., 2025), and provide public access to these reasoning tokens. Although these models are intended to align with “human priors” (Guo et al., 2025), the validity of these reasoning tokens remains a debatable question (Amirizaniani et al., 2024; Dasgupta et al., 2024; Mondorf & Plank, 2024). Testing the validity of thinking traces is beyond the scope of this paper. However, we are interested in a practical objective: leveraging RLMs as a tool to infer latent thinking traces from human experts, thereby improving the reliability of downstream LLM raters.
@@
A powerful paradigm for generating synthetic data is rejection sampling fine-tuning, where a model generates numerous candidate reasoning traces, and only those passing a filtering criterion are used for subsequent training. Existing work has pioneered this approach for tasks with objectively correct answers, such as mathematical reasoning, using an automatic verifier to accept only correct outputs (Guo et al., 2025; Zelikman et al., 2022; Wadhwa et al., 2024; Tong et al., 2024; Singh et al., 2024). To handle more nuanced tasks, methods such as RAFT (Dong et al., 2023) replaced the binary verifier with a learned reward model, filtering for outputs that score highly on desired attributes. This trend of applying reasoning to more diverse domains is also evident in recent work on recommender systems (Tsai et al., 2024), medical question-answering (Chen et al., 2024), software engineering (Cuadron et al., 2024; Penedo et al., 2025), finance (Qian et al., 2025), and cybersecurity (Yu et al., 2025). Concurrently, researchers have begun curating general-purpose SFT datasets enriched with distilled thinking traces to train open reasoning language models (Mattern et al., 2025; Guha et al., 2025). Crucially, while previous work exclusively focuses on using these curated traces for fine-tuning, we introduce a novel second application: distilling insights from the thinking traces to refine annotation codebooks automatically.
@@
The refinement of codebooks is a long-standing challenge in both qualitative research and large-scale data annotation. Research in HCI and crowd-sourcing communities developed human-centric workflows to improve task instructions, often relying on crowd workers to identify ambiguities that were then resolved by an expert or through discussion (Chang et al., 2017; Manam & Quinn, 2018; K. Chaithanya Manam et al., 2019; Pradhan et al., 2022). More recently, researchers have begun to leverage LLMs to automate this process. This includes work on human-LLM collaboration for qualitative coding (Xiao et al., 2023; Halterman & Keith, 2025; Wiebe et al., 2025; Torii et al., 2024; Meng et al., 2025). Other work has demonstrated that LLMs can automatically analyze data or synthesize new guidelines to improve downstream performance (Bibal et al., 2025; Hsu et al., 2025; Srivastava et al., 2025). A key distinction of these automated methods is that they rely only on an LLM’s intrinsic capabilities to analyze content or generate text, rather than being grounded in direct human annotation feedback. In contrast, our approach creates a collaborative pipeline with inputs from both humans and LLMs.
3 Method
3.1 Inferring Latent Thinking Traces
@@
Our goal is to generate a high-quality proxy for this latent trace, which we term the reconstructed thinking trace, $t^{\prime}_{i}$. To achieve this, we use an RLM that can generate thinking tokens, hereafter referred to as the generator model, to perform rejection sampling. For each input item $x_{i}$, we provide the generator model with the codebook $\mathcal{C}$} and the annotation target and prompt it to generate a step-by-step thinking trace that concludes with a final rating. This generation is performed $k$ times independently, yielding a set of $k$ candidate trace-label pairs, $\{(t^{\prime}_{i,j},y^{\prime}_{i,j})\}_{j=1}^{k}$ where $t^{\prime}_{i,j}$ denotes the $j$-th candidate trace and $y^{\prime}_{i,j}$ denotes the $j$-th predicted label for $x_{i}$. For a model to reach the correct human label $y_{i}$, its reasoning is more likely (but not guaranteed) to be a plausible approximation of the latent human thought process. Therefore, we filter the candidate set by retaining only those traces where the model’s predicted label $y^{\prime}_{i,j}$ matches the ground-truth human label $y_{i}$. This creates a set of traces for each sample $x_{i}$:
@@
Figure 2: An example of the inferred thinking trace in evaluating the engagement level of a short story.
Figure 3: Illustration of Inferring Thinking Traces through an RLM. Multiple candidate thinking traces and labels are sampled. Only candidates that are aligned with human-annotated labels are preserved. The detailed process is provided in Section 3.1.
3.2 Improving LLM Raters
@@
$\mathcal{L}_{\text{SFT}}=\mathbb{E}_{(x,t^{\prime},y)\sim\mathcal{D}_{\text{reason}}}\left[-\log\pi_{\theta}(t^{\prime},y|x)\right]$ (1)
@@
We provide the complete prompts for this process in Appendix D.
@@
Step-by-Step Rating
@@
1. The story is very simple, lacks depth, and introduces no intricate or developed elements. 2. The story is somewhat simple, with few elements that are underdeveloped, fragmented, or disconnected. 3. The story demonstrates some complexity, introducing multiple elements, but lacks depth, development, or coherence in integrating these elements. 4. The story is mostly complex, weaving together several developed elements with minor areas that could be expanded or refined. 5. The story is highly elaborate and complex, integrating multiple layers (e.g., detailed world-building, character depth, advanced structure, thematic sophistication) in a cohesive and interconnected way.
@@
Figure 4: An example of the refined annotation codebook. Some content has been omitted due to space limitations. See Appendix E for a complete example.
3.3 Evaluation
@@
This criterion measures the degree to which an LLM rater’s outputs agree with human judgments. We apply this evaluation to the outputs of both our applications. For this analysis, we treat a held-out set of human ratings as the gold standard and compare the LLM-generated labels against them. To quantify this agreement, we use several standard metrics for labels in the Likert scale. Our primary focus is on Kendall’s Tau ($\tau$), as ranking correlation [^3] We observe a high correlation between Spearman’s $\rho$ and Kendall’s $\tau$ in the results. We only report $\tau$ here for simplicity. is the most used measure in the literature (Liu et al., 2023; 2024; Thakur et al., 2025; Gu et al., 2024). For a more complete assessment, we also report the Mean Squared Error (MSE) for average error magnitude and the Intraclass Correlation Coefficient (ICC3) for consistency, based on a two-way mixed-effects model (Koo & Li, 2016).
@@
For codebook refinement, another robust indicator of success is whether $\mathcal{C}$$^{\prime}$ enables independent raters to arrive at the same conclusions consistently. High inter-rater reliability is a sign of a clear and well-defined annotation process. To measure this, we use the improved codebook $\mathcal{C}$$^{\prime}$ to prompt $M$ different proprietary LLMs to rate the same set of items. We then calculate the reliability across these $M$ independent LLM raters using ICC3 (Koo & Li, 2016).
@@
4 Experiment
4.1 Datasets
@@
  3. 3.
    Evaluating News Summaries. We use the SummEval benchmark (Fabbri et al., 2021), which provides human evaluations of machine-generated summaries of news articles. Each summary is annotated by three expert annotators with prior experience in writing research paper summaries for academic conferences. We focus on the consistency and fluency, which respectively evaluate factual correctness and linguistic quality, two key aspects of summary reliability and readability.
  4. 4.
@@
More statistics and pre-processing details for each dataset are provided in Appendix B.
4.2 Data Processing
Figure 5: Trade-off between Data Utilization Rate and Number of Samples. Data utilization rate denotes the percentage of data points that have matched thinking traces within the $k$ generations.
@@ 4.2 Data Processing
As outlined in Section 3, we use rejection sampling to infer thinking traces from the annotation datasets. The original annotation codebook from each source serves directly as the prompt. For our generator models, we selected DeepSeek-R1 and gpt-oss-120b as they output complete thinking traces. We prompt each model to generate sixteen ($k=16$) thinking traces for every sample and keep the first trace that aligns with the human annotator’s final rating. Here, the selection of $k$ is a pragmatic trade-off between computational budget and the coverage of matched valid thinking traces, as shown in Figure 5. A sampling temperature of 1.0 is used during inference to encourage generative diversity.
4.3 Fine-Tuning Specialized LLM Rater
Task Original Model SFT (DeepSeek Trace) SFT (gpt-oss Trace)
Metric $\tau$ ICC3 MSE $\tau$ ICC3 MSE $\tau$ ICC3 MSE
Short Story (Complexity) 0.177 0.231 2.016 0.463* 0.604* 0.838* 0.460* 0.581* 0.944*
Short Story (Engagement) 0.275 0.289 1.939 0.048 0.364 1.364 0.179 0.371 1.831
Student-Written Essay 0.202 0.233 1.947 0.322* 0.267 1.149* 0.240 0.244 1.387*
News Summary (Consistency) 0.141 0.186 2.904 0.276 0.319 3.476 0.243 0.331 2.970
News Summary (Fluency) 0.275 0.378 1.938 0.319 0.402 1.160* 0.306 0.498 1.200*
Translation 0.171 0.240 2.778 0.260* 0.348* 2.492 0.256* 0.314* 2.439*
Chatbot Response (Correctness) 0.162 0.288 2.049 0.297* 0.403* 1.815* 0.263* 0.433* 1.897
Chatbot Response (Helpfulness) 0.176 0.274 1.969 0.264* 0.422* 1.723* 0.295* 0.466* 1.575*
Average 0.197 0.265 2.193 0.281 0.391 1.752 0.280 0.405 1.780
Table 1: Effect of Reasoning-Enhanced SFT on Human-LLM Rater Agreement (DeepSeek-R1-DistillQwen3-8B). For the metrics, higher is better for Kendall’s $\tau$ and ICC3, while lower is better for MSE. An asterisk (*) denotes a statistically significant improvement over the baseline ($p<0.05$).